A charge q1= 300 micro coulomb is at(0,0) and q2= -5/3q1 is at 20 cm mark on the positive x axis . We also have a charge q3= 8/5q2 located at some unknown point on the x axis . Where is we if the net force on q1 is 7.00 N pushing q1 to the left
1
Expert's answer
2021-08-11T14:06:58-0400

First, we have to consider that the two forces exerted on charge 1 since charge 2 and charge 3 are both negative, we have two attractive forces and q3 has to be located then at x<0 because F12 is directed to the left as FT. Then if we analyze the system we have:


"q_1=3\times10^{-4}\,C
\\ q_2=-5\times10^{-4}\,C
\\ q_3=-8\times10^{-4}\,C"


"F_T=-F_{12}+F_{13}=-7\,N"


"F_T=-\dfrac{kq_1q_2}{r^2_{12}}+\dfrac{kq_1q_3}{r^2_{13}}

\\ F_T+\dfrac{kq_1q_2}{r^2_{12}}=\dfrac{kq_1q_3}{r^2_{13}}

\\r_{13} =\sqrt{\dfrac{kq_1q_3}{F_T+\dfrac{kq_1q_2}{r^2_{12}}}}

\\r_{13} =\sqrt{\dfrac{(9\times10^9Nm^2/C^2)(3\times10^{-4}\,C)(-8\times10^{-4}\,C)}{-7\,N+\dfrac{(9\times10^9Nm^2/C^2)(3\times10^{-4}\,C)(-5\times10^{-4}\,C)}{(0.2\,m)^2}}}

\\ r_{13}=0.2529\,m"


The particle is located then at x= - 0.2529 m = - 25.29 cm